OpenStudy (anonymous):

A solution of 2.50g of a compound dissolved in 25.0g of benzene is observed to freeze at 4.30 degree celsius. Calculate the molar mass of the solute. Tf of benzene is 5.5 degree celsius.

OpenStudy (aaronq):

use \(\Delta T = i*m*K_f\) sub \(n=\dfrac{m}{M}\) into \(m=\dfrac{n_{solute}}{kg_{solvent}} \rightarrow m=\dfrac{\dfrac{m}{M}}{kg_{solvent}}\) sub into original: \(\Delta T = i*\dfrac{\dfrac{m}{\color{red}M}}{kg_{solvent}}*K_f\) ; solve for M